OpenStudy (shylovesyou11):

Could someone help? please Michael and Kathryn bowl together and their combined total score for one game was 425 points. Michael’s score was 70 less than twice Kathryn’s. What were their scores? -Kathryn: 275, Michael: 150 - Kathryn: 260, Michael: 165 - Kathryn: 165, Michael: 260 -Kathryn: 150, Michael: 275

OpenStudy (anonymous):

m = 2k - 70 m + k = 425 (2k-70) + k = 425 3k - 70 = 425 3k = 495 k = 165 Katheryn had 165 & Michael had 260. Your welcome :)

OpenStudy (shylovesyou11):

Thank you! =)